Graph the line with -2/3 slope passing through the point -4,-5
Answer by kadash12(6)

First find the line which can be achieved with point slope.
y-(-5)=-2/3(x-(-4))
y+5=-2/3(x+4)
Now distribute the slope.
3(y+5)=-2(x+4)
3y+15=-2x-8
-15 -15 collect like terms
3y=-2x-23
--------- Divide
3
y=-2/3x-23/3
with that you have the function of the graph which points can be placed to graph.
to check the graph place the function here
https://www.desmos.com/calculator